OpenStudy (anonymous):

Question #35: Assessment Progress A city has a population of 380,000 people. Suppose that each year the population grows by 5.75% . What will the population be after 12 years?

OpenStudy (anonymous):

1st year = 380000 * 0.0575 = 21850 + 380000 = 401850 2nd year = 401850 * 0.0575 = add that value to 401850 .. do it till 12th year. you will find your answer.